Bridgton Golf Academy is evaluating new golf practice equipment. The "Dimple-Max" equipment costs $146,000, has a 4-year life, and costs $10,300 per year to operate. The relevant discount rate is 12 percent. Assume that the straight-line depreciation method is used and that the equipment is fully depreciated to zero. Furthermore, assume the equipment has a salvage value of $10,700 at the end of the project’s life. The relevant tax rate is 21 percent. All cash flows occur at the end of the year. What is the equivalent annual cost (EAC) of this equipment?

$45,354.71

$35,978.95

$142,061.59

$46,771.57
Answer #1

Depreciation tax shield = (146,000 / 4) * 21% = 7665

Net operating cash flow after tax = 10300 * ( 1-21%) = -8137

Net cash flow per year = - 8137 + 7665 = -472

After tax salvage value = 10700 * (1-21%) = 8453

Net present value = -146000 - 472 * PVAF ( 12%,4) + 8453 * PVF (12% , 4th year)

= -146000 - 472 * 3.03735+ 8453 * 0.6355

= -142061.7477

EAC = -142061.7477/ 3.03735 = 46771.57 (approx.)

Option , $46,771.57